• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name slifer

    Log entry time 07:56:29 on January13,2002

    Entry number 80224

    keyword=CSR: Owlshift summary

    S.Choi, P.Solvignon, K. Slifer
    Spent few hours to get coda up and running on adaqs1 and start/end of run scripts to load properly.
    Modified high voltages on right arm shower.
    DAQ rate with 5uA on carbon is about 25 Hz.
    Detector checks at 450 MeV with C12 and BeO target
    runs 20182-20190
    1)Cerenkov tube number 5 has many more hits than its neighbors.
    2)Majority of events have 0 track. Doesn't seem reasonable. Software glitch?

    Elastic Data at 476 MeV. 15LH2R. Hydrogen target with radiator.
    runs 20191 -
    1)CODA rate for 1 uA on H2 target w/ radiator is about 55 Hz.
    2) Physics replay: Targ variables show ~20,000 good events while "PID variables"
    show only about 100-200 good events.